# Infinite dSprites for Disentangled Continual Learning: Separating Memory Edits from Generalization

### v2

@@ Abstract
The ability of machine learning systems to learn continually is hindered by catastrophic forgetting, the tendency of neural networks to overwrite existing knowledge when learning a new task. Continual learning methods alleviate this problem through regularization, parameter isolation, or rehearsal, but they are typically evaluated on benchmarks comprising only a handful of tasks. In contrast, humans are able to learn continually in dynamic, open-world environments, effortlessly achieving one-shot memorization of unfamiliar objects and reliably recognizing them under various transformations. To make progress towards closing this gap, we introduce Infinite dSprites, a parsimonious tool for creating continual classification and disentanglement benchmarks of arbitrary length and with full control over generative factors. We show that over a sufficiently long time horizon, the performance of all major types of continual learning methods deteriorates on this simple benchmark. Thus, Infinite dSprites highlights an important aspect of continual learning that has not received enough attention so far: given a finite modelling capacity and an arbitrarily long learning horizon, efficient learning requires memorizing class-specific information and accumulating knowledge about general mechanisms. In a simple setting with direct supervision on the generative factors, we show how learning class-agnostic transformations offers a way to circumvent catastrophic forgetting and improve classification accuracy over time. Our approach sets the stage for continual learning over hundreds of tasks with explicit control over memorization and forgetting, emphasizing open-set classification and one-shot generalization.
@@ 1 Introduction
Continual learning methods are typically evaluated using standard computer vision datasets segmented into disjoint classification tasks that are learned sequentially. Despite advancements in model scale and complexity, even recent continual learning benchmarks like Split-ImageNet-R (Wang et al., 2022a) feature a limited number of tasks and classes. This constraint raises concerns about their ability to accurately reflect open-ended learning scenarios that humans routinely tackle. To address this issue, we introduce Infinite dSprites (idSprites), a continual learning benchmark generator inspired by the dSprites dataset (Matthey et al., 2017). It allows for procedurally generating a virtually infinite progression of two-dimensional shapes in every combination of orientation, scale, and position (see fig. 2). Crucially, by providing the ground truth values of individual factors of variation (FoVs), idSprites paves the way for methods that efficiently exploit the fundamental property of many visual continual classification problems akin to the ones solved by humans, namely the compositional interactions between objects and transformations, such as perspective projections, lighting changes, rigid transformations, deformations etc. In the spirit of the Omniglot Challenge (Lake et al., 2015), and the Abstraction and Reasoning Corpus (Chollet, 2019), we envision solving idSprites as a stepping stone towards continual learning systems that demonstrate human-like intelligence, characterized by sample efficiency, compositional representations, and the ability to perform one-shot generalization and lifelong open-set learning.
@@
We show that all major types of continual learning methods break down on this simple benchmark when tested over a sufficiently long time horizon. This finding indicates that strategies such as regularization, parameter isolation, experience replay, or parameter-efficient adaptation are not sufficient to mitigate catastrophic forgetting in a truly lifelong learning setting. We further demonstrate that even state of the art vision-language models struggle to separate objects from transformations and reliably solve the object re-identification problem via in-context learning. What we need is a new approach that goes beyond preserving, re-learning, or adapting knowledge and instead leverages the compositional nature of continual learning problems by separating task-specific information from universal mechanisms.
@@
Figure 1: Schematic illustration of Disentangled Continual Learning (DCL) with three modules transferred across tasks: (1) an equivariant network that estimates the parameters of the normalization transformation, (2) a normalization module that outputs a canonical representation of the input image, and (3) a buffer that stores class-specific exemplars. During training, we minimize the loss between the output of the normalization module and the corresponding class exemplar. At test time, we return the label of the exemplar closest to the normalized input.
@@
As noted already by McCloskey & Cohen (1989), catastrophic forgetting is caused by destructive model updates, where adjusting model parameters through gradient descent to minimize the cost function of the current task leads to impaired performance on past tasks. Inspired by this observation, we propose a novel paradigm for continual learning centered around the idea of separating memorization from generalization. We hypothesize that catastrophic forgetting can be minimized by decoupling two objectives: accumulating class-specific information that needs to be memorized and learning general mechanisms that transfer well to past and future tasks. We aim to minimize destructive updates by having a separate update procedure for the memory buffer and the generalization model. This separation allows us to expand, maintain, or prune the class-specific knowledge stored in the memory buffer while continually training the generalization model. By learning universal transformations, we can not only avoid destructive gradient updates, but efficiently accumulate knowledge over time. We call this approach Disentangled Continual Learning (DCL)  [^1] Please note the difference from Disentangled Representation Learning.. Section 3.2 describes a proof of concept of DCL in the context of class-incremental continual learning. Our implementation consists of an exemplar buffer that stores a single exemplar per class, an equivariant network that learns to regress the parameters of an affine transform mapping any input into its exemplar, and a normalisation module that applies the predicted affine transformation to the input.
@@
  • We show that continual learning methods are not able to solve a simple benchmark created with idSprites.
  • We propose Disentangled Continual Learning (DCL), a novel approach to continual learning focused on separating explicit memory edits from gradient-based model updates.
  • We implement a proof of concept and show it can efficiently learn over hundreds of tasks with a constant computational budget and a slowly growing memory footprint.
  • We demonstrate how learning data symmetries promotes positive forward and backward transfer, open-set classification, and zero-shot generalisation.
@@ 2 Motivation: Three issues with class-incremental continual learning
Benchmarking
Continual learning datasets are typically limited to just a few tasks and at most a few hundred classes. In contrast, humans can learn and recognise countless novel objects throughout their lifetime. We argue that as the continual learning community, we should focus more on scaling the number of tasks in our benchmarks. We show that when tested over hundreds of tasks standard methods inevitably fail: the effect of regularization decays over time, adding more parameters quickly becomes unpractical, and storing and replaying old examples causes a rapid increase in both memory and compute. Moreover, to tackle individual sub-problems in continual learning, such as the influence of task similarity on forgetting, the role of disentangled representations, and the influence of hard task boundaries, we need to be able to flexibly create datasets that let us isolate these problems. We should also step away from the static training and testing stages and embrace streaming settings where the model can be evaluated an any point.
@@
Continual learning methods are usually benchmarked on class-incremental setups, where a classification problem is split into several tasks to be solved sequentially (van de Ven et al., 2022). Note that the classification learning objective is invariant to identity-preserving transformations of the object of interest, such as rotation, change of lighting, or perspective projection. Unsurprisingly, the most successful discriminative learning architectures, from AlexNet (Krizhevsky et al., 2012) through ResNet (He et al., 2016) to Vision Transformer (Dosovitskiy et al., 2020), learn only features that are relevant to the classification task and discard valuable information about universal transformations, symmetries, and compositionality (Tishby & Zaslavsky, 2015; Higgins et al., 2022). In doing so, they entangle the particular class information with the knowledge about generalization mechanisms and represent both in the weights of the model. When a new task arrives, there is no clear way to update the two separately. We argue that transferring a purely discriminative model across tasks is not conducive to positive forward or backward transfer.
@@
In this paper, we reframe the problem and recognize that the information about identity-preserving transformations, typically discarded, is important for transfer across tasks. For instance, a change in illumination affects objects of various classes similarly. Understanding this mechanism would lead to better generalization on future classes. Consequently, we suggest that modeling these transformations is key to achieving positive forward and backward transfer in continual classification. Symmetry transformations, or equivariances, offer a structured framework for this kind of modeling, which we elaborate on in the subsequent section.
Pre-trained models
@@
Continual adaptation promises to combine the impressive generalization capabilities of large vision-language models trained on extensive swaths of data with the flexibility required for continual learning. While these methods achieve impressive results on many continual learning benchmarks, we argue that their strength comes primarily from the ability of the powerful backbone to solve individual tasks without much adaptation. Zhou et al. (2023) have recently demonstrated that freezing the model and classifying with the average embedding of each class outperforms sophisticated prompt tuning methods such as Learning to Prompt (Wang et al., 2022b) and DualPrompt (Wang et al., 2022a). Similarly, Panos et al. (2023) show that adapting only on the first learning session and then performing Linear Discriminate Analysis on top of a frozen network offers competitive performance on standard benchmarks.
@@
Evaluating new types of methods requires new kinds of benchmarks. As general models pre-trained on Internet-scale data become common, it is important for continual learning benchmarks to still require adapting to never seen data. Considering the emergent zero-shot abilities of foundation models, we should focus less on achieving competitive accuracy on existing benchmarks and more on re-thinking the desiderata, objectives, and metrics we use to design and evaluate continual learning systems. We agree with Chollet (2019) that the measure of intelligence is not only skill, but the speed of skill acquisition. We believe that new continual learning benchmarks should emphasize sample efficiency, positive forward and backward transfer, zero-shot and few-shot generalization, and open-ended learning. We see synthetic data as a perfect tool to measure progress in these areas.
@@ 3 Methods
In this section, we describe in more detail the two main contributions of this work: a software package for generating continual learning benchmarks and a conceptual continual learning framework along with an example implementation. We would like to emphasize that our method serves as a baseline and a proof of concept, showcasing the potential of DCL, and is not intended as a practical method for general use. We see it as a first step towards efficient continual learning methods that match humans in the ability to quickly generalize from very limited data and efficiently decompose any classification problem into particular features that need to be explicitly memorized and general, compositional mechanisms that have to be learned.
@@ 3.1 Infinite dSprites
Infinite dSprites (idSprites) is a software framework designed for easy creation of arbitrarily long continual learning benchmarks. A single idSprites benchmark consists of $T$ tasks, where each task is an $n$-fold classification of procedurally generated shapes. Similar to dSprites, each shape is observed in all possible combinations of the following FoVs: color, scale, orientation, horizontal position, and vertical position. Figure 2 shows an example batch of images with four FoVs and two values per factor (in general, our implementation allows for arbitrary granularity). The canonical form corresponds to a scale of 1, orientation of 0, and horizontal and vertical positions of 0.5. We only use a single color in our experiments for simplicity and to save computation.
@@
The shapes are generated by first randomly sampling the number of vertices from a discrete uniform distribution over a closed integer interval $\llbracket a,b\rrbracket$, then constructing a regular polygon on a unit circle, randomly perturbing the polar coordinates of each vertex, and finally connecting the perturbed vertices with a closed spline of the order randomly chosen from $\{1,3\}$. All shapes are then scaled and centered so that their bounding boxes and their centers of mass align in the canonical form. We also make orientation identifiable by painting one half of the shape black.
@@
The number of tasks, the number of shapes per task, the vertex number interval, the exact FoV ranges, and the parameters of noise distributions for radial and angular coordinates are set by the user, providing the flexibility to control the length, structure, and difficulty of the benchmark. The framework also provides access to the ground truth values of the individual FoVs. idSprites is a pip-installable Python package and hope it will unlock new research directions in continual classification, transfer learning, and continual disentanglement.
3.2 Disentangled learning
@@
As mentioned earlier, in order to efficiently solve idSprites, we need to clearly distinguish between the generalization mechanism that needs to be learned and the class-specific information that has to be memorized. We start by observing that human learning is likely characterized by such separation. Take face recognition, for example. A child is able to memorize the face of its parent but can still get confused by an unexpected transformation, as evidenced by countless online videos of babies failing to recognize their fathers after a shave. Once we learn the typical identity-preserving transformations that a face can undergo, we need only to memorize the particular features of any new face to instantly generalize over many transformations, such as facial expression, lighting, three-dimensional rotation, scale, perspective projection, or a change of hairstyle. Note that while we encounter new faces every day, these transforms remain consistent and affect every face similarly.
@@
Inspired by this observation, we aim to disentangle generalization from memorization by explicitly separating the learning module from the memory buffer in our model design. The memory buffer stores a single exemplar image of each encountered shape. We assume these are given to us by an oracle throughout training, but it would be possible to bootstrap the buffer with a few initial exemplars. The equivariance learning module is a neural network designed to learn the general transformations present in the data.
Our implementation and training objective
We tackle the class-incremental scenario (van de Ven et al., 2022): at each task $n$ we observe a set of data points $D_{n}=\{(x_{n,m},y_{n,m})\}_{m=1}^{N_{n}}$, where $x_{n,m}$ is an image and $y_{n,m}$ is a vector containing the values of the generative factors and the class label. Each data point belongs to one of $C_{n}$ distinct classes. At the start of each task we add the current task exemplars $E_{n}=\{\hat{x}_{i}\}_{i=1}^{C_{n}}$ to the exemplar buffer. We then train a neural network to output a transformation matrix $\hat{\theta}$ that approximates the ground truth affine transformation matrix $\theta$ mapping each input image to its respective exemplar: $\hat{x}=T_{\theta}(x)$. The network is trained continually to minimize the MSE loss between $\hat{\theta}$ and $\theta$. At test time, we use the trained network to regress an affine transformation matrix mapping a previously unseen image to its exemplar. The normalization module applies the predicted transform to the input image using grid sampling with border values for out-of-bound grid locations. The image is then classified through nearest neighbour lookup. Our normalization module is differentiable and therefore allows for backpropagating a loss computed in image space. In practice, we found that this additional reconstruction loss does not improve much over direct FoV supervision.
Discussion

### v3

@@ Abstract
The ability of machine learning systems to learn continually is hindered by catastrophic forgetting, the tendency of neural networks to overwrite previously acquired knowledge when learning a new task. Existing methods mitigate this problem through regularization, parameter isolation, or rehearsal, but they are typically evaluated on benchmarks comprising only a handful of tasks. In contrast, humans are able to learn over long time horizons in dynamic, open-world environments, effortlessly memorizing unfamiliar objects and reliably recognizing them under various transformations. To make progress towards closing this gap, we introduce Infinite dSprites, a parsimonious tool for creating continual classification and disentanglement benchmarks of arbitrary length and with full control over generative factors. We show that over a sufficiently long time horizon, the performance of all major types of continual learning methods deteriorates on this simple benchmark. This result highlights an important and previously overlooked aspect of continual learning: given a finite modelling capacity and an arbitrarily long learning horizon, efficient learning requires memorizing class-specific information and accumulating knowledge about general mechanisms. In a simple setting with direct supervision on the generative factors, we show how learning class-agnostic transformations offers a way to circumvent catastrophic forgetting and improve classification accuracy over time. Our approach sets the stage for continual learning over hundreds of tasks with explicit control over memorization and forgetting, emphasizing open-set classification and one-shot generalization.
@@ 1 Introduction
Continual learning methods are typically evaluated using standard computer vision datasets segmented into disjoint classification tasks. Despite advancements in model scale and complexity, even recent continual learning benchmarks like Split-ImageNet-R (Wang et al., 2022a) feature a limited number of tasks and classes. This constraint raises concerns about their ability to accurately reflect open-ended learning scenarios that humans routinely tackle. To address this issue, we introduce Infinite dSprites (idSprites), a continual learning benchmark generator inspired by the dSprites dataset (Matthey et al., 2017). It procedurally generates a virtually infinite progression of two-dimensional shapes in every combination of orientation, scale, and position (see fig. 2). Crucially, by providing ground truth values of individual factors of variation (FoVs), idSprites paves the way for methods that efficiently exploit a fundamental property of many real-world continual classification problems, namely the compositional interactions between objects and transformations. In the spirit of the Omniglot Challenge (Lake et al., 2015), and the Abstraction and Reasoning Corpus (Chollet, 2019), we envision solving idSprites as a stepping stone towards continual learning systems that demonstrate human-like intelligence, characterized by sample efficiency, compositional representations, and the ability to perform one-shot generalization and lifelong open-set learning.
@@
We show that all major types of continual learning methods eventually break down on this simple benchmark. This finding indicates that strategies such as regularization, parameter isolation, experience replay, or parameter-efficient adaptation are not sufficient to mitigate catastrophic forgetting in a truly lifelong learning setting. We further demonstrate that even state of the art vision-language models struggle to separate objects from transformations and reliably solve the object re-identification problem via in-context learning. What we need is a new approach that goes beyond preserving, re-learning, or adapting knowledge and instead leverages the compositional nature of continual learning problems by separating task-specific information from universal mechanisms.
@@
Figure 1: Schematic illustration of the conceptual Disentangled Continual Learning (DCL) framework with two modules transferred across tasks: 1) the canonicalization network, consisting of an equivariant module that estimates the parameters of a transformation and a normalization module that applies this transformation to the input, and (2) a buffer that stores class-specific exemplars. The canonicalization network is trained continually to map each image to its exemplar. At test time, we return the label of the exemplar closest to the normalized input.
@@
As noted by McCloskey & Cohen (1989), catastrophic forgetting is caused by destructive model updates, where adjusting model parameters through gradient descent to minimize the cost function of the current task impairs performance on past tasks. Inspired by this, we propose a novel paradigm for continual learning centered on separating memorization from generalization. We hypothesize that catastrophic forgetting can be minimized by decoupling two objectives: memorizing class-specific information and learning general mechanisms that transfer well across tasks. We aim to reduce destructive updates by having separate update procedures for the memory buffer and the generalization model. This allows us to expand, maintain, or prune class-specific knowledge in the memory buffer while continuously training the generalization model. By learning universal transformations, we avoid destructive gradient updates and efficiently accumulate knowledge over time. We call this approach Disentangled Continual Learning (DCL)  [^1] Please note the difference from Disentangled Representation Learning.. Section 3.2 describes a simple implementation of DCL. It consists of an exemplar buffer that stores a single exemplar per class, an equivariant network that learns to regress parameters of an affine transform mapping any input to its exemplar, and a normalization module that applies the predicted affine transformation to the input.
Contributions
@@
  • We show that all major continual learning methods break down on a simple benchmark created with idSprites.
  • We propose Disentangled Continual Learning (DCL), a novel approach to continual learning based on separating explicit memory edits from gradient-based model updates.
  • We implement a proof of concept of DCL and demonstrate it can efficiently learn over hundreds of tasks and perform open-set classification and zero-shot generalization.
@@ 2 Motivation: Three issues with class-incremental continual learning
2.1 Benchmarking
Continual learning datasets are typically limited to just a few tasks and at most a few hundred classes. In contrast, humans can learn and recognise countless novel objects throughout their lifetime. We argue that we should focus more on scaling the number of tasks in our benchmarks. We show that when tested over hundreds of tasks, standard methods inevitably fail: the impact of regularization decays over time, adding more parameters quickly becomes impractical, and replaying old samples eventually becomes ineffective under a constant computational budget. Moreover, to tackle individual sub-problems in continual learning, such as the influence of task similarity on forgetting, the role of disentangled representations, and the impact of hard task boundaries, we need to flexibly create datasets that allow us to isolate these issues. We should also move away from static training and testing stages and embrace streaming settings where the model can be evaluated at any point.
@@ 2.2 Invariant representations
Continual learning methods are usually benchmarked on class-incremental setups, where a classification problem is split into several tasks to be solved sequentially (van de Ven et al., 2022). Note that the classification learning objective is invariant to identity-preserving transformations of the object of interest, such as rigid transformations, deformations, change of lighting, or perspective projection. Unsurprisingly, the most successful discriminative learning architectures, from AlexNet (Krizhevsky et al., 2012) through ResNet (He et al., 2016) to Vision Transformer (Dosovitskiy et al., 2020), learn only features relevant to the classification task and discard valuable information about universal transformations, symmetries, and compositions (Tishby & Zaslavsky, 2015; Higgins et al., 2022). By doing so, they entangle specific class information with knowledge about generalization mechanisms and represent both in the model’s weights. When a new task arrives, there is no clear way to update these separately. We argue that transferring a purely discriminative model across tasks is not conducive to positive forward or backward transfer.
@@
In this paper, we reframe the problem by recognizing that information about identity-preserving transformations, typically discarded, is crucial for transfer across tasks. For instance, changes in illumination affect objects of various classes similarly. Understanding this mechanism can lead to better generalization on future classes. Consequently, we propose that modeling these transformations is key to achieving positive forward and backward transfer in continual classification. Symmetry transformations, or equivariances, provide a structured framework for this modeling, which we elaborate on in the subsequent section.
2.3 Pre-trained models
@@
Continual adaptation promises to combine the impressive generalization capabilities of large vision-language models trained on extensive amounts of data with the flexibility required for continual learning. While these methods achieve impressive results on many continual learning benchmarks, we argue that their strength comes primarily from their powerful backbone and its ability to solve individual tasks without much adaptation. Zhou et al. (2023) recently demonstrated that freezing the model and classifying with the average embedding of each class outperforms sophisticated prompt tuning methods such as Learning to Prompt (Wang et al., 2022b) and DualPrompt (Wang et al., 2022a). Similarly, Panos et al. (2023) show that performing adaptation in the first learning session followed by Linear Discriminate Analysis on top of a frozen network offers competitive performance on standard benchmarks.
@@
Evaluating new types of methods requires new benchmarks. As general models pre-trained on Internet-scale data become common, it is important for continual learning benchmarks to remain challenging, requiring genuine adaptation and generalization to previously unseen data. Considering the emergent zero-shot abilities of foundation models, we should focus less on achieving competitive accuracy on existing benchmarks and more on re-thinking the desiderata, objectives, and metrics we use to design and evaluate continual learning systems. We agree with Chollet (2019) that the measure of intelligence is not only skill, but the speed of skill acquisition. We believe that new continual learning benchmarks should emphasize sample efficiency, positive forward and backward transfer, zero-shot and few-shot generalization, and open-ended learning. We see synthetic data as a perfect tool to measure progress in these areas.
@@ 3 Methods
In this section, we describe in more detail the two main contributions of this work: a software package for generating continual learning benchmarks and a conceptual continual learning framework along with an example implementation. We would like to emphasize that our approach serves as a baseline and a proof of concept, showcasing the potential of DCL, and is not intended as a practical method for general use. We see it as a first step towards efficient continual learning techniques that match humans in the ability to quickly generalize from very limited data and efficiently decompose any classification problem into specific features that need to be explicitly memorized and general mechanisms that need to be learned.
@@ 3.1 Infinite dSprites
Infinite dSprites (idSprites) is a software framework designed for the easy creation of arbitrarily long continual learning benchmarks. A single idSprites benchmark consists of $T$ tasks, where each task is an $n$-fold classification of procedurally generated shapes. Similar to dSprites, each shape is observed in all possible combinations of the following FoVs: color, scale, orientation, horizontal position, and vertical position. Figure 2 shows an example batch of images with four FoVs and two values per factor (in general, our implementation allows for arbitrary granularity). The canonical form corresponds to a scale of 1, an orientation of 0, and horizontal and vertical positions of 0.5. We only use a single color in our experiments for simplicity and to save computation.
@@
The shapes are generated by first randomly sampling the number of vertices from a discrete uniform distribution over a closed integer interval $\llbracket a,b\rrbracket$, then constructing a regular polygon on a unit circle, randomly perturbing the polar coordinates of each vertex, and finally connecting the perturbed vertices with a closed spline of an order randomly chosen from $\{1,3\}$. All shapes are then scaled and centered so that their bounding boxes and their centers of mass align in the canonical form. We also make orientation identifiable by painting one half of the shape black.
@@
The number of tasks, the number of shapes per task, the vertex number interval, the exact FoV ranges, and the parameters of noise distributions for radial and angular coordinates are set by the user, providing the flexibility to control the length, structure, and difficulty of the benchmark. The framework also provides access to the ground truth values of the individual FoVs. idSprites is a pip-installable Python package that we hope will unlock new research directions in continual classification, transfer learning, and continual disentanglement.
3.2 Disentangled continual learning
@@
As mentioned earlier, in order to efficiently solve idSprites, we need to clearly distinguish between the generalization mechanism that needs to be learned and the class-specific information that has to be memorized. We start by observing that human learning is likely characterized by such separation. Take face recognition, for example. A child can memorize the face of its parent but may still get confused by an unexpected transformation, as evidenced by countless online videos of babies failing to recognize their fathers after a shave. Once we learn the typical identity-preserving transformations that a face can undergo, we only need to memorize the particular features of any new face to instantly generalize over many transformations, such as facial expressions, lighting, three-dimensional rotation, scale, and perspective projection. Note that while we encounter new faces every day, these transforms remain consistent and affect every face similarly. Indeed, this fundamental property of the physical world makes generalization possible. As Taylor et al. (2021) aptly note, “to successfully generalize information for appropriate use in novel situations, learned information should reflect stable features of the environment and discard idiosyncratic details of individual experiences.”
@@
Inspired by this observation, we aim to disentangle generalization from memorization by explicitly separating the learning module from the memory buffer in our model design. The memory buffer stores a single exemplar image of each encountered shape. We assume these are provided by an oracle throughout training, but it would be possible to bootstrap the buffer with a few initial exemplars. The equivariance learning module is a neural network designed to encapsulate the general transformations present in the data by learning to canonicalize each input, i.e. transform it to the canonical form. At test time, each input image is canonicalized and then compared to the stored exemplars. This approach draws inspiration from prototype and exemplar-based models of categorization in neuroscience, which operate on the premise that the brain compares the current stimulus with representations of all pertinent categories and selects the one perceived as most similar (Bowman & Zeithamova, 2018; Nosofsky & Johansen, 2000).
3.2.1 Implementation and training objective
At each task, we observe a training set $D_{n}$ of triplets, each comprising an image $x$, its generative factors $y$, and a class label $c$. Since we are tackling the class-incremental scenario (van de Ven et al., 2022), we do not have access to task labels. However, we do assume that for each class we are given an exemplar: a single image showing the shape in the canonical form. Whenever a new class is encountered, we add its exemplar to the memory buffer. We continually train a neural network to regress the parameters of a two-dimensional affine transformation that maps each image to its class exemplar. We supervise the network directly with MSE loss on the transformation parameters, since we can access the generative factors for each input image and calculate the ground truth transform. At test time, we use this network to canonicalize images of previously unseen shapes and compare them with exemplars stored in the buffer. Each image is then classified as belonging to the class of its nearest exemplar. Please see algorithm 1 and algorithm 2 for details of the implementation.
3.2.2 Discussion